# Revisiting Constant Negative Rewards for Goal-Reaching Tasks in Robot Learning

### v1

@@ Abstract
Many real-world robot learning problems, such as pick-and-place or arriving at a destination, can be seen as a problem of reaching a goal state as soon as possible. These problems, when formulated as episodic reinforcement learning tasks, can easily be specified to align well with our intended goal: $-1$ reward every time step with termination upon reaching the goal state, called minimum-time tasks. Despite this simplicity, such formulations are often overlooked in favor of dense rewards due to their perceived difficulty and lack of informativeness. Our studies contrast the two reward paradigms, revealing that the minimum-time task specification not only facilitates learning higher-quality policies but can also surpass dense-reward-based policies on their own performance metrics. Crucially, we also identify the goal-hit rate of the initial policy as a robust early indicator for learning success in such sparse feedback settings. Finally, using four distinct real-robotic platforms, we show that it is possible to learn pixel-based policies from scratch within two to three hours using constant negative rewards. Our video demo can be found here. [^1] Video demo: https://drive.google.com/file/d/1O8D3oCWq5xf2hi1JOlMBbs6W1ClrvUFb/view?usp=sharing
@@ 1 Introduction
In Reinforcement Learning (RL), the task designer implicitly specifies the desired behavior using the reward signal. In order to guide the learning agent to a reasonable solution, task designers often rely on task-specific domain knowledge to hand-craft a dense reward function with state-to-state differences that facilitate faster learning. Since guiding rewards reflect the task designer’s preferred behaviors, they could bias the solutions that the agent finds, potentially leading to sub-optimal outcomes (Riedmiller et al., 2018). For most goal-reaching problems, there is a simpler alternative that is easy to specify but still incorporates our intended goal accurately: a constant negative reward every time step, with termination upon reaching the goal state. We call this the minimum-time specification since maximizing the undiscounted sum of these rewards leads to reaching the goal state as soon as possible.
Many real-world tasks can naturally be specified using sparse rewards, where the agent receives a positive discounted reward only upon reaching the goal state. Minimum-time specifications are a special case of sparse rewards, with the advantage of being simpler due to their undiscounted nature. By using the minimum-time formulation, the task designer can avoid biasing the final solution and focus solely on recognizing task success rather than guiding the learned policy. This approach is also suitable for specifying unforeseen tasks where a human instructor may be unable to provide domain knowledge beforehand.
@@ 2 Related Work
Minimum time problems in optimal control aim to transfer a system from an initial to a final state in the shortest time, with the cost function being the time taken between steps (Chui and Chen, 2012, Chapter 9). For example, Penicka et al., (2022) combines classical path planning with model-free deep RL to optimize a neural network policy for minimum-time flight of a quadrotor through a sequence of waypoints with obstacles. Control benchmarks in RL like MountainCar (Sutton and Barto, 2018) and DotReacher (Garg et al., 2022) also use the minimum-time formulation.
@@ 3 How to Specify Goal-Reaching Tasks in RL?
Guiding reward formulation
The reward function is inspired by Brockman et al., (2016), where reacher is incentivized to get as close to the target as possible. We also include a commonly used precision reward term previously used on a real robot reacher task using the UR5 industrial arm (Lan et al., 2022; Farrahi and Mahmood, 2023; Che et al., 2023; Elsayed et al., 2024). The reward function has two components: (i) penalty term calculated as the negative of the Euclidean distance between the fingertip and the target, and (ii) precision term where the agent receives $+1$ reward when the fingertip is inside the target (see Table 1). Each episode has a time limit of 1000. The time limit is set long enough to incentivize the agent to remain at the goal once reached.
Contact Reward Formulation
The agent will receive a reward of $+1$ whenever it is within the target sphere and 00 otherwise (see Table 1). Each episode has a time limit of $1000$. This choice of reward and timeout follows the dm_control suite (Tassa et al., 2018).
Minimum-Time formulation
Shorter episodes are encouraged by imposing a reward of $-1$ for each step until termination. The episode terminates when the fingertip of the arm reaches the target with near-zero velocity. An episode is completed only when the agent reaches the goal state. If the agent exceeds the time limit, we reset the environment such that the goal state remains the same, but the robot arm moves to a different starting state. Note that, unlike the previous formulations, we get rid of the fixed time limit horizon, where all episodes have a uniform episode length.
@@ 5 Does Choice of Time Limit Impact Success in Minimum-Time Tasks?
6 Learning Vision-Based Policies on Minimum-Time Robot Tasks
@@ Acknowledgments
We would like to thank Richard Sutton, Matthew Taylor, Patrick Pilarski, Shivam Garg, Varshini Prakash and Bram Grooten for several insightful discussions that helped improve the quality and clarity of the paper. We are also appreciative of the computing resources provided by the Digital Research Alliance of Canada and the financial support from the CCAI Chairs program, the RLAI laboratory, Amii, and NSERC of Canada. We are also thankful to the Ocado Group and Huawei Noah’s Ark Lab for their generous donation of the UR5 and Franka Emika Panda, respectively.

### v2

# Revisiting Sparse Rewards for Goal-Reaching Reinforcement Learning

@@ Abstract
Many real-world robot learning problems, such as pick-and-place or arriving at a destination, can be seen as a problem of reaching a goal state as soon as possible. These problems, when formulated as episodic reinforcement learning tasks, can easily be specified to align well with our intended goal: $-1$ reward every time step with termination upon reaching the goal state, called minimum-time tasks. Despite this simplicity, such formulations are often overlooked in favor of dense rewards due to their perceived difficulty and lack of informativeness. Our studies contrast the two reward paradigms, revealing that the minimum-time task specification not only facilitates learning higher-quality policies but can also surpass dense-reward-based policies on their own performance metrics. Crucially, we also identify the goal-hit rate of the initial policy as a robust early indicator for learning success in such sparse feedback settings. Finally, using four distinct real-robotic platforms, we show that it is possible to learn pixel-based policies from scratch within two to three hours using constant negative rewards. Our video demo can be found here. [^1] Video demo: https://youtu.be/a6zlVUuKzBc
@@ 1 Introduction
In Reinforcement Learning (RL), the task designer implicitly specifies the desired behavior using the reward signal. In order to guide the learning agent to a reasonable solution, task designers often rely on task-specific domain knowledge to hand-craft a dense reward function with state-to-state differences that facilitate faster learning. Since guiding rewards reflect the task designer’s preferred behaviors, they could bias the solutions that the agent finds, potentially leading to sub-optimal outcomes (Riedmiller et al., 2018). For most goal-reaching problems, there is a simpler alternative that is easy to specify but still incorporates our intended goal accurately: a constant negative reward every time step, with termination upon reaching the goal state. We call this the minimum-time specification since maximizing the undiscounted sum of these rewards leads to reaching the goal state as soon as possible. This specification avoids biasing the final solution, focusing solely on recognizing task success. This approach is also suitable for specifying unforeseen goal-reaching tasks where a human instructor may be unable to provide domain knowledge beforehand.
An alternative and simple specification of rewards is to give a discounted positive reward upon “success” and zero everywhere else. Such a reward is often considered a sparse reward, as a differentiating reward signal is available only sparsely. The minimum-time reward can be seen as an extreme form of sparse reward, where the reward value never changes and the differentiating signal only comes from episodic accumulated rewards, called the return. However, minimum-time specification has the advantage of being simpler as it does not require discounting as part of the problem.
@@ 2 Related Work
Minimum time problems in optimal control aim to transfer a system from an initial to a final state in the shortest time, with the cost function being the time taken between steps (Chui and Chen, 2012, Chapter 9). For example, Penicka et al., (2022) combines classical path planning with model-free deep RL to optimize a neural network policy for minimum-time flight of a quadrotor through a sequence of waypoints with obstacles. Control benchmarks in RL like MountainCar (Sutton and Barto, 2018) and DotReacher (Garg et al., 2022) also use the minimum-time formulation with a reward of $-1$ each step until termination.
@@ 3 How to Specify Goal-Reaching Tasks in RL?
Guiding reward formulation The reward function is inspired by Brockman et al., (2016), where reacher is incentivized to get as close to the target as possible. We also include a commonly used precision reward term previously used on a real robot reacher task using the UR5 industrial arm (Lan et al., 2022; Farrahi and Mahmood, 2023; Che et al., 2023; Elsayed et al., 2024). The reward function has two components: (i) penalty term calculated as the negative of the Euclidean distance between the fingertip and the target, and (ii) precision term where the agent receives $+1$ reward when the fingertip is inside the target (see Table 1). Each episode has a time limit of 1000. The time limit is set long enough to incentivize the agent to remain at the goal once reached.
Contact Reward Formulation The agent will receive a reward of $+1$ whenever it is within the target sphere and 00 otherwise (see Table 1). Each episode has a time limit of $1000$. This choice of reward and timeout follows the dm_control suite (Tassa et al., 2018).
Minimum-Time formulation Shorter episodes are encouraged by imposing a reward of $-1$ for each step until termination. The episode terminates when the fingertip of the arm reaches the target with near-zero velocity. An episode is completed only when the agent reaches the goal state. If the agent exceeds the time limit, we reset the environment such that the goal state remains the same, but the robot arm moves to a different starting state. Note that, unlike the previous formulations, we get rid of the fixed time limit horizon, where all episodes have a uniform episode length.
@@ 5 Does Choice of Time Limit Impact Success in Minimum-Time Tasks?
6 Learning Vision-Based Minimum-Time Tasks on Real Robots
@@ Acknowledgments
We would like to thank Richard Sutton, Matthew Taylor, Patrick Pilarski, Shivam Garg, Varshini Prakash and Bram Grooten for several insightful discussions that helped improve the quality and clarity of the paper. We would also like to thank Homayoon Farrahi for providing the training script that helped us learn the UR5-Reacher task using a curriculum. We are also appreciative of the computing resources provided by the Digital Research Alliance of Canada and the financial support from the CCAI Chairs program, the RLAI laboratory, Amii, and NSERC of Canada. We are also thankful to Ocado Technology and Huawei Noah’s Ark Lab for their generous donation of the UR5 and Franka Emika Panda, respectively.